OpenStudy (anonymous):

If m<8 =22, m<7=31 and m

OpenStudy (anonymous):

∠7 + ∠8 +∠9 = ∠ead 31+22+∠9 = 102 53 + ∠9 = 102 ∠9 = 102-53 ∠9=49